# A Tale of Single-channel Electroencephalogram: Devices, Datasets, Signal Processing, Applications, and Future Directions

### v1

@@ Abstract
Single-channel electroencephalogram (EEG) is a cost-effective, comfortable, and non-invasive method for monitoring brain activity, widely adopted by researchers, consumers, and clinicians. The increasing number and proportion of articles on single-channel EEG underscore its growing potential. This paper provides a comprehensive review of single-channel EEG, focusing on development trends, devices, datasets, signal processing methods, recent applications, and future directions. Definitions of bipolar and unipolar configurations in single-channel EEG are clarified to guide future advancements. Applications mainly span sleep staging, emotion recognition, educational research, and clinical diagnosis. Ongoing advancements of single-channel EEG in AI-based EEG generation techniques suggest potential parity or superiority over multichannel EEG performance.
@@
EEG, Single-channel, Wearable Devices, EEG Processing, Sleep Staging, Emotion Recognition, Education Assessment, Clinical Diagnosis, EEG Signal Generation.
@@ I INTRODUCTION
Electroencephalography (EEG) is a technique for recording the brain’s electrical signals, which has high temporal resolution and non-invasiveness, making it an indispensable tool in brain research and neurological disorder diagnosis [1]. EEG measures the electrical activity by detecting voltage fluctuations resulting from ionic current flows within the neurons of the brain [2]. By recording electrical activity from the scalp, EEG can provide the real-time insights into brain function, enabling the investigation of cognitive processes under different conditions and mechanisms of neurological disorders [3]. Traditionally, EEG signals are taken on the surface of the scalp, while the intracranial EEG (iEEG) directly records electrical signals inside the brain using implanted electrodes [4]. In this paper, we focus primarily on traditional scalp EEG signals, in which the International 10-20 system, shown in Fig1, is commonly used for EEG electrode placement and for correlating the external skull locations to the underlying cortical areas [5].
@@
In clinical settings, EEG is essential for diagnosing and monitoring conditions such as epilepsy and sleep disorders, as well as assessing brain function during cognitive tasks, aiding in the understanding of neural mechanisms underlying perception, memory, and decision-making [6, 7, 8, 9, 10]. Moreover, EEG is widely applied in non-clinical settings due to its portability and cost-effectiveness, which facilitates monitoring of emotions and stress for mental health management, tracks students’ attention in education to enhance the teaching process, and aids in optimizing concentration and relaxation during sports and meditation training for effective outcomes [11, 12, 13]. However, traditional EEG systems often employ multiple channels to capture a detailed spatial map of brain activity. For instance, high-density EEG systems can improve the spatial resolution but present some challenges such as usability, convenience, and high cost [14]. These limitations pose significant barriers to the widespread application of EEG beyond controlled lab settings [15]. The complexities involved in setup and maintenance hinder the use of high-density EEG systems in clinical, consumer, industrial sectors, and other situations. In addition, wearable electrode caps can cause physical discomfort, limiting recording duration and frequency, particularly in clinical, consumer, and industrial sectors, and in studies involving children or the elderly.
@@
Single-channel EEG is a simple system that applies two electrodes to record single-channel electrical activity, which can be classified into unipolar single-channel EEG and bipolar single-channel EEG based on the electrode placement and reference selection [16], as shown Fig2. Bipolar single-channel EEG means that two electrodes are placed at different locations on the scalp and the potential difference between these two electrodes reflects the difference in electrical activity between two specific brain areas. The unipolar reference is used when all the electrodes are referenced to a unique physical reference or a unique virtual reference [17]. The physical reference is typically an electrode (e.g., Cz, Fz and Oz) placed on the scalp or body surface during the recording setup. The virtual reference is a linear combination of recordings from all electrodes, usually obtained during offline processing after EEG data acquisition. Typical examples of virtual references include the reference electrode standardization technique (REST), average reference (AR), and linked-mastoids/ears reference (LM) [18, 19, 20]. The primary difference is that unipolar single-channel EEG employs one recording electrode and one reference electrode, whereas bipolar single-channel EEG uses two recording electrodes to measure the relative potential changes between them. Single-channel EEG systems can be divided into two types: wired and wireless. Wired single-channel EEG systems are mostly part of a multi-channel system, while wireless single-channel systems are mostly from wireless portable devices.
Single-channel EEG is effectively used in sleep studies to monitor sleep stages and detect disorders like sleep apnea [21, 22]. It can integrate into wearable devices for continuous home monitoring, which is convenient and less intrusive for patients [23]. In stress monitoring, single-channel EEG detects changes in brainwave patterns linked to stress and relaxation, offering feedback for stress reduction interventions [24]. Similarly, in cognitive load assessment, single-channel EEG measures brain activity associated with mental effort, beneficial in educational settings and user experience research [25].
As a simplified EEG recording technique, single-channel EEG shows significant potential in clinical and non-clinical applications. However, there remains a lack of systematic reviews on single-channel EEG, encompassing both unipolar and bipolar configurations. Furthermore, the definition of single-channel EEG is not strictly standardized, posing challenges in scientific research. Notably, the issue of optimal electrode positioning for specific tasks has not been adequately addressed. This review aims to summarize existing research on single-channel EEG, discussing its devices, datasets, signal processing, applications, challenges, and future development directions.
In Section II, we review the evolution of EEG and explore research trends in single-channel EEG. Section III introduces the design and implementation of consumer-grade single-channel EEG devices. Section IV covers commonly used datasets for single-channel EEG research. Section V discusses signal processing methods. Section VI explores single-channel EEG applications in tasks such as sleep staging, disease diagnosis, and emotion recognition. Section VII summarizes and discusses the content of the preceding sections, highlighting the main challenges and future development directions in single-channel EEG research. Section VIII concludes this review, aiming to promote further development and application of single-channel EEG technology.
@@ II TRENDS OF SINGLE-CHANNEL EEG
A literature survey was conducted to highlight trends in single-channel EEG research over the past 5 years. The total number of articles containing the keyword ”Single-channel EEG” from 2019 to 2023 was retrieved from Google Scholar. All studies related to EEG were identified using the keyword ”EEG”. Non-relevant articles, where the search keyword appeared only as part of the author’s name, were excluded. The survey revealed an increasing focus on single-channel EEG studies between 2019 and 2023, as depicted in Figure3. The number of publications on single-channel EEG rose from 841 in 2019 to 1,050 in 2021, reaching 1,750 in 2023. The ratio of Single-channel EEG to general EEG studies increased from 0.94% in 2019 to 2.01% (2021) and 2.77% (2023). This represents a nearly six-fold increase in the proportion of single-channel EEG studies over the period, reflecting growing recognition and attention from researchers. Furthermore, advancements in signal processing and data analysis methods for single-channel EEG have significantly contributed to the rapid development of this field [31].
Figure 3: Total number of general and single-channel EEG publications over the past 5 years, as searched on Google Scholar. Single-channel EEG is an increasingly popular research topic. (Publications up to June 2024).
@@ III DEVICES of SINGLE-CHANNEL EEG
The NeuroSky MindWave series and Sichiray EEG devices are the typical single-channel EEG headsets, each with different features. For instance, the NeuroSky MindWave is an affordable single-channel EEG device designed for education, entertainment, and research, which uses the ThinkGear ASIC Module (TGAM) [32] module to measure and output EEG power spectra such as $\alpha$, $\beta$ waves, etc., providing an easy-to-use brainwave monitoring solution. This device consists of a headset with a T-shaped headband, wide ear clips, and flexible sensor arms that rest on the forehead above the eyes, which connects via bluetooth low energy and classic bluetooth, with a single AAA battery providing approximately 8 hours of use [33]. The NeuroSky MindWave Mobile offers similar functionality to the original MindWave but with improvements in comfort and connection [34]. Its headsets feature automatic wireless pairing, enhanced ergonomics, and work seamlessly with a wide range of educational and gaming apps available on the NeuroSky App Store. The MindWave Mobile II uses a TGAM2.9B Bluetooth v4.0 module and features a flexible rubber sensor arm, rounded forehead sensor tip, T-shaped headband, and wide ear clips for a more ergonomic design and a more comfortable wearing experience. It also uses a single AAA battery with a battery life of 8 hours and is compatible with Windows, Mac, iOS, and Android platforms [35]. The OpenBCI EEG headset kit, especially the Ultracortex Mark IV, is a more advanced and customizable option for professional research that requires high-precision and multi-channel EEG data. It offers a high degree of customization options suitable for single-channel EEG applications [36].
The Sichiray 2.0 is another single-channel EEG device designed for consumer and educational use [37], which offers similar functionality to the NeuroSky MindWave series but provides fewer details about its specific features, connection options and supported applications.
TableI presents a summary of the single-channel EEG devices used in academic literature, including information about their release dates, available sensors and data availability. The market for single-channel EEG devices is currently limited, with major brands including the NeuroSky MindWave series, Sichiray 2.0, and OpenBCI devices, which are primarily used in education, entertainment and preliminary research. Future advancements in technology and growing application needs are expected to enhance these devices in terms of accuracy, comfort, endurance and versatility. This will enable them to better meet the demands of professional research and practical applications.
@@
  Device Released Sensors Data Available????
  NeuroSky MindWave[33] 2011 Fp1 Y
NeuroSky MindWave Mobile [34] 2012 Fp1 Y
Olimex EEG-SMT [38] 2013 User-configurable Y
NeuroSky MindWave Mobile II [35] 2018 Fp1 Y
OpenBCI EEG Headband Kit [36] 2022 User-configurable Y
Sichiray [37] 2022 Fp1 Y

### v2

@@ Abstract
Single-channel electroencephalogram (EEG) is a cost-effective, comfortable, and non-invasive method for monitoring brain activity, widely adopted by researchers, consumers, and clinicians. The increasing number and proportion of articles on single-channel EEG underscore its growing potential. This paper provides a comprehensive review of single-channel EEG, focusing on development trends, devices, datasets, signal processing methods, recent applications, and future directions. Definitions of bipolar and unipolar configurations in single-channel EEG are clarified to guide future advancements. Applications mainly span sleep staging, emotion recognition, neurofeedback, educational research, and clinical diagnosis. Additionally, we discuss about the artificial intelligence (AI)-based EEG generation techniques, advancements through the integration of advanced signal processing with AI, innovations in hardware development, and strategies for the integration of wearables enabled by the Internet of Things (IoT), collectively establishing a foundational roadmap for future developments in single-channel EEG systems and their applications.
@@
EEG, Single-channel, Wearable Devices, EEG Processing, Sleep Staging, Emotion Recognition, Education Assessment, Clinical Diagnosis, Neurofeedback, EEG Signal Generation, Hardware, IoT.
@@ I INTRODUCTION
Electroencephalography (EEG) is a technique for recording the brain’s electrical signals, which has high temporal resolution and non-invasiveness, making it an indispensable tool in brain research and neurological disorder diagnosis [1]. EEG measures the electrical activity by detecting voltage fluctuations resulting from ionic current flows within the neurons of the brain [2]. By recording electrical activity from the scalp, EEG can provide the real-time insights into brain function, enabling the investigation of cognitive processes under different conditions and mechanisms of neurological disorders [3]. Traditionally, EEG signals are taken on the surface of the scalp, while the intracranial EEG (iEEG) directly records electrical signals inside the brain using implanted electrodes [4]. In this paper, we focus primarily on traditional scalp EEG signals, in which the International 10-20 system, shown in Fig. 1, is commonly used for EEG electrode placement and for correlating the external skull locations to the underlying cortical areas [5].
@@
In clinical settings, EEG is essential for diagnosing and monitoring conditions such as epilepsy and sleep disorders, as well as assessing brain function during cognitive tasks, aiding in the understanding of neural mechanisms underlying perception, memory, and decision-making [6, 7, 8, 9, 10]. Moreover, EEG applications extend to non-clinical settings through its portability and cost-effectiveness, enabling emotion and stress monitoring for mental health management, attention tracking in educational environments, and concentration optimization in sports and meditation training [11, 12, 13]. However, traditional EEG systems often employ multiple channels to capture a detailed spatial map of brain activity. High-density EEG systems enhance spatial resolution while introducing challenges in usability, convenience, and cost efficiency [14]. Multi-channel EEG systems necessitate extensive skin preparation and adhesive conductive gels, resulting in time-consuming processes and user discomfort that impede repeated measurements [15, 16]. Conductive gel application introduces signal quality variability and setup complexities in non-clinical environments [17], while electrode caps induce physical discomfort, constraining recording duration and applicability for children and elderly subjects [18]. These limitations, combined with setup and maintenance complexities, pose significant barriers to widespread EEG adoption beyond laboratory settings [19]. In contrast, single-channel EEG, particularly with dry electrodes [20, 21], reduces setup complexity and enhances user comfort, thus making EEG more accessible for broader applications.
@@
Single-channel EEG is a simple system that applies two electrodes to record single-channel electrical activity, which can be classified into unipolar single-channel EEG and bipolar single-channel EEG based on the electrode placement and reference selection [22], as shown Fig. 2. Bipolar single-channel EEG means that two electrodes are placed at different locations on the scalp and the potential difference between these two electrodes reflects the difference in electrical activity between two specific brain areas. The unipolar reference is used when all the electrodes are referenced to a unique physical reference or a unique virtual reference [23]. The physical reference is typically an electrode (e.g., Cz, Fz and Oz) placed on the scalp or body surface during the recording setup. The virtual reference is a linear combination of recordings from all electrodes, usually obtained during offline processing after EEG data acquisition. Typical examples of virtual references include the reference electrode standardization technique (REST), average reference (AR), and linked-mastoids/ears reference (LM) [24, 25, 26]. The primary difference is that unipolar single-channel EEG employs one recording electrode and one reference electrode, whereas bipolar single-channel EEG uses two recording electrodes to measure the relative potential changes between them. Single-channel EEG systems comprise wired variants, predominantly integrated within multi-channel systems, and wireless variants, primarily implemented in portable devices. Single-channel EEG enables critical health monitoring applications, particularly in sleep stage analysis and sleep apnea detection, while facilitating non-intrusive home monitoring through wearable device integration [27, 28, 29]. Beyond sleep studies, this technology detects brainwave pattern changes associated with stress and relaxation states for stress reduction interventions, and measures brain activity linked to mental effort for cognitive assessment in educational settings and user experience research [30, 31].
@@
As a simplified EEG recording technique, single-channel EEG shows significant potential in clinical and non-clinical applications [33]. Balam et al.[34] systematically reviewed single-channel EEG-based drowsiness detection methods, addressing safety implications and analytical techniques while excluding broader single-channel EEG applications. An early mini-review [35] explored wireless and wearable EEG systems and brain-computer interfaces, primarily focusing on broader technological advancements rather than single-channel applications. A specialized study examining ocular artifact denoising in single-channel EEG signals advanced signal quality enhancement [36], yet maintained a narrow focus excluding broader applications and configurations. Despite the comprehensive analysis of multi-channel systems [37], the unique characteristics of single-channel EEG system configurations remain relatively unexplored. Therefore, a narrative review is essential for summarizing research on single-channel EEG, highlighting its devices, datasets, applications, challenges, and future development directions, thereby addressing the growing importance of this technology.
A literature search was conducted in the Google Scholar database up to October 2024 using the search terms “Single-channel” AND “EEG” OR “electroencephalography” OR “devices”; “Single-channel” AND “EEG” OR “electroencephalography” OR “datasets”; “Single-channel” AND “EEG” OR “electroencephalography” OR “signal processing”; “Single-channel” AND “EEG” OR “electroencephalography” OR “applications”. Additionally, we searched for “single-channel EEG” AND “sleep staging” OR “emotion recognition” OR “education research” OR “clinical diagnosis” OR “EEG signal generation” OR “bipolar” OR “unipolar” OR “neurofeedback” OR “AI” OR “deep learning” OR “generative model” OR “hardware” OR “IoT”. We also retrieved relevant articles cited in other narrative reviews and screened references from original papers to ensure no significant reports were overlooked. We included articles that specifically discussed single-channel EEG in terms of its technical development, usage in different contexts, advancements in signal processing, and future potential.
In Section II, we review the evolution of EEG and explore research trends in single-channel EEG. Section III addresses consumer-grade single-channel EEG device design and implementation. Section IV presents datasets utilized in single-channel EEG research. Section V analyzes signal processing methodologies. Section VI investigates single-channel EEG applications encompassing sleep staging, disease diagnosis, emotion recognition, neurofeedback, and educational research. Section VII evaluates the preceding content while examining challenges and future development trajectories in single-channel EEG research. Section VIII concludes by addressing implications for single-channel EEG technology advancement.
@@ II TRENDS OF SINGLE-CHANNEL EEG
Although our literature search encompassed publications up to October 2024, the incomplete volume of articles for 2024 was taken into account by focusing our trend analysis on studies from 2019 to 2023. The total number of articles containing the keyword “Single-channel EEG” from 2019 to 2023 was retrieved from Google Scholar. All studies related to EEG were identified using the keyword “EEG”. Non-relevant articles, where the search keyword appeared only as part of the author’s name, were excluded. The survey revealed an increasing focus on single-channel EEG studies between 2019 and 2023, as depicted in Fig. 3. The number of publications on single-channel EEG rose from 841 in 2019 to 1,050 in 2021, reaching 1,750 in 2023. The ratio of Single-channel EEG to general EEG studies increased from 0.94% in 2019 to 2.01% (2021) and 2.77% (2023). This represents a nearly six-fold increase in the proportion of single-channel EEG studies over the period, reflecting growing recognition and attention from researchers. Furthermore, advancements in signal processing and data analysis methods for single-channel EEG have significantly contributed to the rapid development of this field [42].
Figure 3: Total number of general and single-channel EEG publications over the past 5 years, as searched on Google Scholar. Single-channel EEG is an increasingly popular research topic.
@@ III DEVICES of SINGLE-CHANNEL EEG
This section introduces the commonly used single-channel EEG devices in detail, focusing on their instrumentation and measurement characteristics, including features such as sensor, analog-to-digital converter (ADC) resolution, sampling rate, connectivity, battery life, max power consumption and data availability.
@@
  Device Released Sensor ADC Resolution (bits) Sampling Rate (Hz) Electrode Type Connection Price (USD) Battery Life (Hours) Max Power Consumption Data Available?
  NeuroSky MindSet [43] 2007 Fp1 12 512 Dry Bluetooth 199.99 - 15mA @ 3.3V Y
NeuroSky Mindflex [44] 2009 Fp1 12 512 Dry Wired 79.99 - 15mA @ 3.3V N
NeuroSky MindWave [45] 2011 Fp1 12 512 Dry Bluetooth $9$ 15mA @ 3.3V Y
XWave headset [46] 2011 Fp1 12 512 Dry Wired 99 - 15mA @ 3.3V N
NeuroSky MindWave Mobile [47] 2012 Fp1 12 512 Dry Bluetooth 129.99 - 15mA @ 3.3V Y
NecoMimi [48] 2012 Fp1 12 512 Dry Bluetooth 69 - 15mA @ 3.3V N
Olimex EEG-SMT [49] 2013 Fpz or Oz 10 256 Dry - - 3 - Y
MyndPlay Mindband [48] 2014 Fp1 12 512 Dry BLE 219 10 15mA @ 3.3V Y
Aurora Dreamband [50] 2015 Fp1 - - Dry BLE 299 8 15mA @ 3.3V Y
NeuroSky MindWave Mobile II [51] 2018 Fp1 12 512 Dry Bluetooth/BLE 129.99 - 15mA @ 3.3V Y
Macrotellect Brainlink lite V2 [52] - Fp1 12 512 Dry Bluetooth 179 5 15mA @ 3.3V Y
Macrotellect Brainlink Pro [52] - Fp1 12 512 Dry BLE 259 4 15mA @ 3.3V Y
Sichiray 2.0 [53] 2022 Fp1 12 215 Dry/Wet BLE 80 8 5mA @ 3.3V N
  • BLE: Bluetooth Low Energy, a power-efficient version of Bluetooth. AAA: Standard AAA batteries are used to power the device. Y: Data is available for access or download. N: Raw data is unavailable for access or download. -: No relevant information found.
EEG devices utilize two main types of electrodes: wet and dry [20]. Wet electrodes require conductive gel to reduce impedance at the skin-electrode interface, ensuring high-quality signal acquisition and are considered the gold standard [16]. However, the need for gel increases preparation time and limits device portability. In contrast, dry electrodes are made from materials such as conductive polymers or metals that can capture electrical signals from the scalp without the need for conductive gel [21]. This design makes dry electrodes more convenient and user-friendly, particularly for single-channel EEG devices. Additionally, semi-dry electrodes represent a compromise between wet and dry types, using minimal electrolyte to improve signal quality while maintaining relative convenience. Among these options, dry electrodes are widely used as they simplify setup, enhance portability and suit wearable and long-term monitoring.
NeuroSky was one of the pioneering companies to produce and distribute consumer-grade EEG devices in the market. In 2007, NeuroSky launched their first product, the MindSet headset, which utilizes a dual-electrode system with an ear-mounted reference and forehead-positioned signal capture electrode [43]. Building on this initial success, NeuroSky expanded their product line with more advanced devices like the MindWave, a cost-effective single-channel EEG solution that incorporates the ThinkGear ASIC Module (TGAM) [54], a specialized chip featuring 12-bit ADC resolution, 512Hz sampling rate and 15mA @ 3.3V power consumption, to measure and output EEG power spectra such as $\alpha$, $\beta$ waves and derived metrics (attention, meditation), providing an easy-to-use brainwave monitoring solution. Additionally, The NeuroSky MindWave consists of a T-shaped headband with wide ear clips and forehead sensors above the eyes, connects via Bluetooth, and operates for 8 hours using a single AAA battery [45]. The NeuroSky MindWave Mobile improves upon the original with better comfort and connectivity, featuring automatic wireless pairing, enhanced ergonomics, and compatibility with various educational and gaming apps from their App Store [47]. The MindWave Mobile II further enhances comfort through its TGAM2.9B Bluetooth v4.0 module, flexible rubber sensor arm, rounded forehead sensor tip, T-shaped headband and wide ear clips, while offering 8-hour battery life from a single AAA battery and compatibility with Windows, Mac, iOS and Android platforms [51]. For developers interested in custom implementations, NeuroSky provides the TGAM module at \$49 per piece, offering comprehensive measurement capabilities including blink detection, mental effort, familiarity, appreciation, emotional spectrum, creativity, and alertness measurement [55]. This core technology has been widely implemented, powering both NeuroSky’s own products (MindSet [43], MindWave series [47], Mindflex [44]) and third-party devices (XWave headset [46], NecoMimi [48], MyndPlay Mindband [48], Macrotellect Brainlink [52]). While based on the same TGAM architecture, these devices are optimized for distinct application domains, ranging from cognitive research and educational assessment to interactive entertainment systems. For instance, Mindflex [44] enables users to control a foam ball through obstacles using brainwaves, while NecoMimi [48] features plush cat ears that respond to brainwave activity. In a similar market segment, the Sichiray 2.0 provides consumer and educational EEG monitoring with a compact design, operating at 3.3V @ 5mA via Bluetooth communication. The device incorporates a micro USB charging port, 3-125 Hz frequency range, 12-bit ADC, and 215 Hz sampling rate, outputting eSense metrics including $\delta$ and $\theta$ waves for mental state analysis while utilizing an onboard notch filter for 50 Hz interference elimination [53].
Technological advancements have driven single-channel device proliferation, leading manufacturers to develop diverse consumer-oriented solutions. MyndPlay’s TGAM-based headband integrates recreational features with cognitive development functionalities [48], while the Aurora Dreamband [50] incorporates sensor technology for movement and sleep cycle monitoring. These consumer-grade devices facilitate cognitive state research without extensive technical expertise [21] while expanding into education and gaming applications [56]. Their economic accessibility enhances user adoption while maintaining core functionalities. Despite performance constraints in attention and meditation studies, environmental calibration integration, multi-sensor fusion, and deep learning technologies have enhanced their educational adaptability. NeuroSky devices demonstrate educational research efficacy despite simplified signal acquisition [13]. Compared to multi-channel systems such as Emotiv [57], Muse [58], and OpenBCI [59], single-channel devices offer enhanced convenience for entry-level research requiring moderate precision. The BrainLink Lite V2 and Pro [52] implementations utilize single-channel (Fp1) capabilities for raw EEG acquisition and eSense-based monitoring. The Pro variant features specialized earclips, enhanced relaxation protocols, cardiac monitoring, EmoLight functionalities, and BLE 4.0 connectivity, surpassing Lite’s Bluetooth 3.0. Pro additionally incorporates PPG heart rate sensors and expanded BCI capabilities, offering comprehensive accessory packages versus Lite’s basic configuration. Additionally, the Pro model integrates PPG heart rate sensors and expanded BCI capabilities, with both variants offering mobile app compatibility, though Pro provides a more comprehensive accessory package versus Lite’s basic headband configuration. Additionally, The Olimex EEG-SMT utilizes a 256 Hz sampling rate and two-channel differential input 10-bit ADC, though operationally employing a single channel configuration with dual active electrodes (CH- and CH+) for voltage differential measurement [49].
In the field of instrumentation and measurement, single-channel devices will maintain significance as algorithms and sensor technology advance, with future developments focusing on intelligent signal processing optimization and enhanced device adaptability and accuracy. User experience and usability factors, including wearing comfort for long-term use, wireless connectivity for simplified data transmission, and compact design for improved portability, are critical considerations that influence their widespread application. Table I summarizes the key characteristics of single-channel EEG devices documented in academic literature, covering specifications from released time and sensor type to ADC resolution, sampling rate, electrode configuration, connectivity options, pricing, battery duration, and data accessibility. These devices exhibit varying technical capabilities, particularly in their ADC resolution and sampling rates, while their choice of connection method (Bluetooth or wired) further impacts the quality of signal capture and efficiency of data transmission - factors whose importance varies depending on specific research needs. Looking forward, comprehensive user studies focusing on comfort, operational simplicity, and overall satisfaction will be essential to guide future design optimizations and enhance the user experience.